Obtain the equation of the conic, a focus of which lies at (2,1), the directrix of

which is x+y = 0 and which passes through (1,4). Also identify the conic.


1
Expert's answer
2021-05-21T07:38:15-0400

Given,

Focus of the conic is lies at "(2,1)"

The directrix of the conic lies at ",x+y=0"

It passes through the point "(1,4)"

Distance between the given points "(d)=\sqrt{1+9}=10"

The distance from the point "(1,4)" from directrix "=\frac{5}{\sqrt{2}}"

The ratio of distance "=\frac{\sqrt{20}}{5}"

From the above, we can conclude that the ratio is less than 1, so it is an ellipse.

Hence the required equation is "\frac{(x-2)^2+(y-1)^2}{(\frac{(x+y)}{\sqrt{2}})^2}=(\frac{2}{\sqrt{5}})^2"

"\Rightarrow \frac{x^2+4-4x+y^2+1-2y}{\frac{x^2+y^2+2xy}{2}}=\frac{4}{5}"

"\Rightarrow 5x^2+20-20x+5y^2+5-10y=2x^2+2y^2+4xy"

"\Rightarrow 3x^2-4xy+3y^2-20x-10y+25=0"